If the length of rhombus is 13 cm and the one diagonal is 24 cm . what is the area of the rhombus
Answer by rothauserc(4718)

We know that the diagonals of a rhombus bisect each other at right angles, then
:
(24/2)^2 + x^2 = 13^2
:
x^2 = 169 - 144 = 25
:
x = 5
:
therefore the other diagonal is 2(5) = 10
:
******************************************
Area of rhombus = (10 * 24) / 2 = 120 cm^2
******************************************
: